OpenStudy (anonymous):

Evaluate the following indefinite integral Just to double check my answer @ganeshie8 @hartnn

OpenStudy (anonymous):

\[\int\limits_{}^{}x(2-\frac{ 1 }{ x^3 })dx\]

OpenStudy (anonymous):

my final answer is\[x^2+\frac{ 1 }{ x }+c\]

OpenStudy (anonymous):

the easiest way to check if your answer is correct is to differentiate it.

OpenStudy (anonymous):

But how to differnetiate c @Beauregard

OpenStudy (anonymous):

differentiate*

OpenStudy (anonymous):

What is the derivative of a constant/

OpenStudy (anonymous):

err, i'm not sure

OpenStudy (dan815):

its 0 u silly billy

OpenStudy (dan815):

if u were to graph y=3

OpenStudy (dan815):